OpenStudy (anonymous):

Which expression can be used to determine the nth term in the pattern below? 21, 30, 39, 48, 57, ...

OpenStudy (_michaeld22):

30-21 = 9. 39 - 30 = 9. It seems like you're adding 9 each time.. Soo, 57+9 = ?

OpenStudy (anonymous):

add 9 4 times

OpenStudy (solomonzelman):

\(\Large\color{black}{ \bf a_n=a_{n-1}+9 }\)

OpenStudy (uri):

There is always a formula for questions like these @_Michaeld22 No use of teaching them directly. :)

OpenStudy (_michaeld22):

Touche.. Lol I don't remember the formulas tho, I just try to help the best I can, and get people like @SolomonZelman to type them for me (:

OpenStudy (solomonzelman):

Yeah, I am not in the mood of teaching recursive formulas for this type of sequences. I assume that people have a proper background to the problem they are doing. And if they don't, I apologize for the inconvenience, and leave the question.